Electron. J. M ath. Phys. Sci. 2002 , 1 , 1 ,00–00

Electronic Journal of Mathematical and Physical Sciences

EJMAPS

ISSN: 1538-263X www.ejmaps.org

TIME VARIABLE BAYESIAN SAMPLE TEST PLAN: A MATH-EMATICAL MODEL FOR A REAL TIME FAULT TOLERANT COMPUTER SIMULATION PLATFORM

Abraham F. Jalbout^{1,2}

1 Department of Chemistry,University of New Orleans, New Orleans, LA 70148-2820 2 Department of Physics and Engineering, Dillard University, 2601 Gentilly Blvd., New Orleans, LA 70112

Abstract: This paper is focused on thedevelopment of a Bayesian sample test plan for an industrial setting. Insoftware engineering the testing of real time (RT) or near real time (NRT)fault tolerant software can be applied to almost any practical situation. The developmentof such platform can lead to a ground braking industrial shift in currentmethodology. The reason, being that the Bayesian class of methods provide forquicker and more convergent methods of statistical analysis. A completemathematical model for a new computational platform based on a rigorousmathematical model is presented.

Keywords: Bayesian,fraction defective, computer platform, real time default system

\begin{matrix}c \\ ◯ \end{matrix}2002 by EJMAPS. Reproduction for noncommercial purposes permitted.